### v1

# CloudFix: Automated Policy Repair for Cloud Access Control Policies Using Large Language Models

Abstract
Access control policies are vital for securing modern cloud computing, where organizations must manage access to sensitive data across thousands of users in distributed system settings. Cloud administrators typically write and update policies manually, which can be an error-prone and time-consuming process and can potentially lead to security vulnerabilities. Existing approaches based on symbolic analysis have demonstrated success in automated debugging and repairing access control policies; however, their generalizability is limited in the context of cloud-based access control. Conversely, Large Language Models (LLMs) have been utilized for automated program repair; however, their applicability to repairing cloud access control policies remains unexplored. In this work, we introduce CloudFix, the first automated policy repair framework for cloud access control that combines formal methods with LLMs. Given an access control policy and a specification of allowed and denied access requests, CloudFix employs Formal Methods-based Fault Localization to identify faulty statements in the policy and leverages LLMs to generate potential repairs, which are then verified using SMT solvers. To evaluate CloudFix, we curated a dataset of 282 real-world AWS access control policies extracted from forum posts and augmented them with synthetically generated request sets based on real scenarios. Our experimental results show that CloudFix improves repair accuracy over a Baseline implementation across varying request sizes. Our work is the first to leverage LLMs for policy repair, showcasing the effectiveness of LLMs for access control and enabling efficient and automated repair of cloud access control policies. We make our tool Cloudfix and AWS dataset publicly available.
I Introduction
Cloud compute systems allow administrators to write policies that specify access control rules, which specify how access to private data is governed. Policies are typically written in a convenient language, such as the Amazon Web Services (AWS) Identity and Access Management (IAM) policy language, allowing administrators to write their own policies. However, this is a tedious, complex, and often error-prone process. Moreover, policies must be modified on a regular basis due to the ever changing nature of cloud computing. Creating and maintaining policies is a challenging endeavor and often requires significant technical skills and domain expertise in ensuring access control is correctly implemented. This can potentially lead to smaller organizations or those without the proper resources to forgo best security practices in ensuring access privilege is kept to a minimum, and instead provide broad access in order to ensure their cloud computing services work as intended.
Cloud computing services often provide tools to assist administrators in writing and maintaining policies or monitoring access to resources (e.g., AWs IAM Access Analyzer, AWS IAM Policy Simulator, AWS Cloudtrail service). However, when policies do not function as they are intended, or must be modified to incorporate access to new services or resources, these tools are inadequate when administrators lack the technical expertise to use them. This is additionally compounded by the lack of automated approaches for creating or repairing policies, as existing techniques either do not extend to cloud access control policies or are restricted in the ways they repair policies [Son2013FixMU, peng2017towards, D'Antoni2024, DetecingResolvingPolicyMisconfig, Jayaraman2014AutomatedAA]. What is needed is automated techniques that can reason about the complexities of cloud policies and be able to modify and repair existing policies as permissions and services evolve throughout the cloud ecosystem.
Recently, Large Language Models (LLMs) have seen great success in the field of software engineering [LLM4SESurvey]. They have become invaluable tools for software developers and engineers with their ability to generate code and automate tasks. In particular, LLMs have been shown to be useful for automated program repair, where LLMs generate patches to fix faulty programs [zhang2024systematic, li2025hybrid, hossain2024deep, xia2023automated, campos2025empirical, orvalho2025counterexample]. In the field of access control, LLMs have been shown to be successful in synthesizing policies from natural language descriptions [jayasundara2023sokaccesscontrolpolicy, vatsa2025synthesizingaccesscontrolpolicies, 1269406, EnforcingAccessControlLLMS]. However, their applicability in the context of repairing access control policies has yet to be explored.
In this work, we take the first step towards leveraging Large Language Models for access control policy repair. Given a faulty or misconfigured access control policy and access control requests which must be allowed or denied by the policy, our approach repairs the policy so that it correctly classifies the given set of requests. We implement our approach in a tool called CloudFix and showcase its effectiveness in repairing AWS IAM access control policies. Additionally, we curate the first policy repair dataset for cloud access control policies and make it publicly available for future research endeavors. Specifically, we make the following key contributions:
  • We automate the access control policy repair process using an LLM-driven Fault-localization-guided approach.
  • We present the first integration of LLMs with SMT-based formal verification for repairing access control policies;
  • We experimentally show that our fault-localization techniques enable LLMs to repair faulty policies.
  • We release the public dataset for cloud policy repair and make the source code for CloudFix available on GitHub.
The paper is organized as follows. In Section II we provide the overview access control policies and the policy repair problem. In Section III we introduce CloudFix. In Section IV we describe how we curate a dataset to evaluate CloudFix. In Section V we experimentally validate CloudFix. In Section VI we discuss our findings. In Section VII we discuss the related work. In Section VIII we conclude our paper.
II Overview and Background
In this work we focus primarily on access control policies written in AWS IAM policy language. However, the techniques we present in this work are not specific to the AWS IAM policy language and can be applied to any policy language.
II-A Amazon Web Service Policies
An AWS IAM policy $P$ is a declarative policy consisting of a list of statements describing how access is governed. Policies field access control requests which are either allowed or denied by the policy. An access control request is a tuple $(\delta,a,r,e)$ where $\delta$ is a specific principal performing the access, $a$ is the action to be performed, $r$ is the resource being accessed, and $e$ is the set of environment variables within the request (e.g., IP address, current time, etc). A policy statement $S$ is a 5-tuple (Effect, Principal, Action, Resource, Condition) where:
  • Effect is either Allow or Deny
  • Principal consists of a list of users or entities
  • Action is a list of actions
  • Resource is a list of resources
  • Condition is an optional list of conditions further specifying how access is constrained
Initially, all requests are implicitly denied. An access control request is granted access if and only there is at least one statement within the policy which explicitly allows access and no statement in the policy explicitly revokes or denies access. The AWS policy language allows special characters to be used which represent broad access: the ’*’ wildcard character can be used to represent any string, while the ’?’ character can be used to represent any character. Additionally AWS allows identity-based policies which omit the Principal field.
Figure 1: Manual Policy Repair Process. Stage 1 shows an administrator authoring a policy while a user generates access requests. In Stage 2, a cloud evaluation engine validates these requests, correctly allowing one (s3:GetObject) but implicitly denying another (ec2:RunInstances) because no statement permits it. In Stage 3, the administrator manually interprets this feedback, adding a new Allow statement to the policy to fix the issue, and resubmits it for validation.
II-B A Motivating Example
In real-world cloud computing, debugging and repairing access control policies are done collaboratively but largely manually. Administrators write policies in the AWS IAM consoles and attach them to users/roles, or use the IAM Policy Simulator, which still requires manual work. This prompts iterative manual debugging until the desired permissions are allowed and denied. A common yet risky solution for this is to craft overly broad permissions (e.g., action: s3: *) to temporarily prevent requests from being unintentionally denied and attempt to narrow the scope iteratively.
Figure 1 demonstrates how an AWS engineer identifies errors in the policy and attempts to repair it. The process has three states: in Stage 1, the Cloud Security Admin writes an access control policy that gives access to the principal to read any s3 category and object across the account. In the first request, the user is asking to read the object newfile.txt stored in the s3 category my-category by invoking the s3:GetObject action. In the second request, the user is asking to launch new ec2 instances and access the corresponding backend logs in the us-west-2 region by performing the ec2:RunInstances action on the resource arn:aws:ec2:us-west-2:backend/logs. These requests are stored in CloudTrail as access history logs and are accessible to the Admin. Next, the Admin examines the requests through the Cloud Engine to locate the failure points in the policy, as shown in Stage 2. The Cloud Engine validates and returns that the first request is correctly allowed, as the policy has an explicit statement that allows it, and the second request is implicitly denied since there is no statement allowing the request. In Stage 3, the Admin adds an explicit statement in the policy that allows the required request, which was initially denied by the policy. The Admin then uses the Cloud Engine to validate the repaired policy and deploys it.
This manual repair and validation process is tedious, time-consuming, and error-prone. Each time a user’s request is denied or overly allowed, the administrator must inspect the IAM policy line by line, trace the evaluation logic through the Cloud Engine, and cross-check the results against CloudTrail logs to understand why the policy denied or allowed the access request. Manually inspecting the policy to repair it requires redeploying and retesting the policy to confirm that intended requests are still allowed and denied. Moreover, this iterative process heavily relies on human reasoning and repeated trial and error; it becomes increasingly impractical as the number of users, policies, and requests grows.
III CloudFix
In this section we introduce our policy repair framework CloudFix. The framework consists of four steps: (A) Goal Validation, (B) Fault Localization, (C) Prompt Generation, and (D) Repair Synthesizer. Figure 2 demonstrates the overall stages of the method. CloudFix operates over an input access control policy and request specification. The Goal Validator first evaluates the policy against the request specification: if all requests are correctly classified, the policy is considered correct and the framework returns the policy. Otherwise, the process proceeds iteratively. In each iteration, the faulty policy and the requests are fed into the Fault Localizer component that analyzes the failed requests and the current policy to identify statements that are responsible for misclassifications. These faulty statements, together with the policy and requests, are provided to the Prompt Generator, which produces an input prompt for the Repair Synthesizer. The Repair Synthesizer queries an LLM to generate a candidate policy as a potential repair. This candidate policy is validated against the requests by the Goal Validator. If the candidate policy correctly classifies all requests, then the policy has been completely repaired, and the algorithm terminates. Otherwise, the candidate policy becomes the new input to the next iteration of the loop. This process continues until a fully repaired policy is produced, or the iteration threshold is reached. Algorithm 1 presents the complete CloudFix algorithm.
III-A Goal Validator
The Goal Validator determines if the policy correctly classifies all requests in the given specification. Algorithm 2 shows how a request specification, $r$ in a set $R$ is validated against a policy $P$ to test whether $r$ is not allowed by $P$ and determine that $P$ is faulty. To do this, both the policy $P$ and $r$ are encoded into logical formulas suitable for satisfiability checking using the Quacky tool [eiers2022quacky]. The algorithm returns FAIL if at least one request is misclassified. This information is then passed to the Fault Localizer for further identification of statements within $P$ that are responsible for the misclassification of the $r$ in $S$. Encoding the requests and the policies as SMT formulas using an SMT Solver enables us to eliminate the dependency on external policy evaluation engines. We use a maximum of $I$ iterations per policy to reach a repair goal of 100%. This can be defined by the user depending on the available resources. Once the target accuracy or the maximum $I$ is reached, the repair algorithm stops and returns the repaired policy. If the target repair accuracy is not achieved within $I$ iterations, the algorithm retains the repair with the highest accuracy.
Algorithm 1 RepairPolicy
1:Policy $P$, Requests $R=\{R_{\text{allow}},R_{\text{deny}}\}$, Maximum iterations II
2:Repaired Policy $P^{*}$ or Failure
[…]
4:if $verification=\textsc{pass}$ then
5:  return $P$ $\triangleright$ Policy already satisfies requests
6:else
7:  for $i=1$ to II do
8:   $S_{f}\leftarrow\textsc{LocalizeFaults}(P,R)$
9:   $prompt\leftarrow\textsc{GeneratePrompt}(P,R,S_{f})$
10:   $P^{*}\leftarrow\textsc{SynthesisRepairWithLLM}(prompt)$
11:   $\mathit{verification}\leftarrow\textsc{ValidateGoal}(P^{*},R)$
12:   if $\mathit{verification}=\textsc{pass}$ then
13:     return $P$
14:   else if P* gets better than P then
15:     $P=P^{*}$
16:   end if
17:  end for
18:  return Failure
19:end if
Algorithm 2 ValidateGoal
1:Policy $P$, Requests $R=\{R_{\text{allow}},R_{\text{deny}}\}$
2:PASS or FAIL indicating whether $P$ is faulty
[…]
4:for each $(\delta,a,r,e)\in R_{\text{allow}}$ in $R$ do
5:  if $(\delta,a,r,e)\not\models\Phi$ then
6:   return FAIL $\triangleright$ Missed allow
7:  end if
8:end for
9:for each $(\delta,a,r,e)\in R_{\text{deny}}$ inRR do
10:  if $(\delta,a,r,e)\models\Phi$ then
11:   return FAIL $\triangleright$ Missed deny
12:  end if
13:end for
14:return PASS $\triangleright$ Policy satisfies all goals
III-B Fault Localizer
Once the Goal Validator determines that a policy $\mathit{P}$ is faulty (i.e., it misclassifies requests), then $\mathit{P}$ is passed to the Fault Localizer to identify statements within it that misclassify the requests. We define a request misclassification as allowing a request that is supposed to be denied and/or denying a request that is supposed to be allowed by the policy. We define three types of faults within the policy that are responsible for misclassifications: case 1) faulty statements that explicitly allow requests which should be denied as per the request specification, case 2) statements that explicitly deny requests that should be allowed per the request specification, and case 3) requests which are implicitly denied yet should be allowed. The process is shown in Algorithm 3. Each misclassified request is analyzed with respect to the policy to determine the type of fault that incorrectly classifies it.
Given a policy $\mathit{P}$ and a request set $\mathit{R}$, lines 4 - 17 in Algorithm 3 illustrate how we identify faulty statements within $\mathit{P}$. The $\mathrm{SMTSolver}$ function determines the satisfiability of policy constraints against specific requests using the Z3 solver [z3] for satisfiability checking and the Quacky tool for policy and request encoding. For each misclassified request, to check if it falls into case 1, $\mathrm{SMTSolver}(r,s)$ iteratively checks each allow statement $\mathit{s\in S^{+}}$ to see if it allows request $\mathit{r}$ by creating a minimal policy containing only that statement. Lines 7 and 8 show that if $\mathrm{SMTSolver}(r,s)$ returns $\mathit{SAT}$, we assert that there is an explicit allow statement that is allowing the statement that should have been denied. To check case 2, $\mathrm{SMTSolver}(r,P^{\star}\cup\{s\}))$ combines the universal policy with a deny statement, $\mathit{s}$, to determine if the deny statement explicitly denies $\mathit{r}$. This is shown in lines 11-12. If the solver returns $\mathit{UNSAT}$, it indicates that the deny statement in $\mathit{S}$ is responsible for the incorrect denial of $\mathit{r}$. If this condition fails, case 3 holds, as shown in line 14. Finally, Algorithm 3 returns a list that has a mapping between policy statements and request failures, along with the failure types. This combination of statements, along with the explicit requests they allow, deny, or implicitly deny, is then passed to the Prompt Generator, which is discussed below.
Figure 2: This flowchart details the architecture of the CloudFix framework.
III-C Prompt Generator
Once the Fault Localizer identifies the faulty statements $s$ within $P$, we format the results into a structured natural language prompt for the LLM. The Prompt Generator receives the fault localization report, the list of request specifications $R$, and the policy $P$. It then structures them into a format suitable for LLMs. We annotate the prompts with the iteration number and the accuracy obtained from the Goal Validator to provide contextual feedback to the LLM. In this work, we consider two types of prompts: Base and Fault Localization-Guided.
Base Prompt
This is the simplest prompt we employ to simulate a real-world scenario where an engineer has access only to a policy and requests. It consists of the policy that needs to be repaired in accordance with the must-allow and must-deny list specifications.
Fault Localization Guided Prompt
We construct this prompt by first identifying the requests that each statement in a policy misclassifies. This augments the base prompt by providing the LLM with detailed information that maps specific failures to their root causes within the policy using an SMT-based fault localization report. Figure 3 illustrates an example. The prompt includes the policy, the must-allow and must-deny requests, and a fault-localization finding that indicates the error type, the failed requests, and the responsible statements for the failure.
POLICY:
Statement:{[
{Sid: VisualEditor0, Effect: Allow, Action: iam:GetAccountSummary, Resource: *}
{Sid: VisualEditor2, Effect: Allow, Action: s3:*, Resource: *}
{ Sid: VisualEditor3, Effect: Deny, Action: ec2:DescribeInstances, Resource: * }]}
REQUESTS:
Effect: Deny,
Action: [s3:GetObject],
Resource: [arn:aws:s3:::admin-category/document.txt]
Effect: Allow,
Action: [sqs:SendMessage],
Resource: [arn:aws:athena:us-east-1:123456789012:workgroup/primary]
Effect: Deny,
Action: [sqs:SendMessage],
Resource: [arn:aws:s3:::admin-category/document.txt]
Effect: Deny,
Action: [ec2:DescribeInstances],
Resource:[arn:aws:glue:us-east-1:123456789012:table/my-database/my-table]
Effect: Allow,
Action: [ec2:DescribeInstances],
Resource:[arn:aws:athena:us-east-1:123456789012:workgroup/primary]
FAULT LOCALIZATION:
SHOULD BE DENIED BUT EXPLICITLY ALLOWED
Action: [s3:GetObject]
Resource: [arn:aws:s3:::admin-category/document.txt]
Expected: deny, Got: allow
Responsible stmt: VisualEditor2
SHOULD BE ALLOWED BUT IMPLICITLY DENIED
Action: [sqs:SendMessage]
Resource: [arn:aws:athena:us-east-1:123456789012:workgroup/primary]
Expected: allow, Got: deny
Responsible stmt: None (missing allow)
SHOULD BE ALLOWED BUT EXPLICITLY DENIED
Action: [ec2:DescribeInstances]
Resource: [arn:aws:athena:us-east-1:123456789012:workgroup/primary]
Expected: allow, Got: deny
Responsible stmt: VisualEditor3
Figure 3: An example of a structured input provided to the LLM for a fault-localization-guided repair. The prompt is composed of three parts: the original faulty POLICY, the set of REQUESTS, and the FAULT LOCALIZATION report. The report explicitly maps failed requests to the policy statements responsible for the errors, categorizing them into three types of failures: "denied but explicitly allowed", "allowed but implicitly denied", or "allowed but explicitly denied".
III-D Repair Synthesizer (LLM)
We leverage an LLM to generate candidate repairs directly from misclassified requests and fault localization findings. The Repair Synthesizer LLM produces a candidate policy, which is then passed to the Goal Validator for verification. If the Goal Validator determines that the policy correctly classified the request specification, the policy will be returned as a successful repair. However, due to the complexity of the requests and the policies themselves, some faulty policies are difficult to fully repair in a single iteration. The repair cycle continues until the Goal Validator determines that the candidate policy correctly classifies the given must-allow and must-deny request specifications, or the predefined iteration threshold is reached. The candidate policy which achieved the highest classification accuracy is returned as a partially repaired policy.
Algorithm 3 LocalizeFaults
1:Misclassified requests $R$, policy $P$ with statements $S$
2:Fault list $L$
[…]
5:Let $P^{\star}$ be the universal-allow policy
[…]
7:for each $r\in R$ do
8:  if $r.\mathrm{expected}=\mathrm{Deny}$ and $r.\mathrm{actual}=\mathrm{Allow}$ then
9:   $A_{r}\leftarrow\{\,s\in S^{+}\mid\mathrm{SMTSolver}(r,\{s\})=\mathrm{SAT}\,\}$
10:   $L\leftarrow L\cup\{(r,A_{r})\}$ $\triangleright$ explicit allow
11:  else if $r.\mathrm{expected}=\mathrm{Allow}$ and $r.\mathrm{actual}=\mathrm{Deny}$ then
12:   $D_{r}\leftarrow\{\,s\in S^{-}\mid\mathrm{SMTSolver}(r,P^{\star}\cup\{s\})=\mathrm{UNSAT}\,\}$
13:   if $D_{r}\neq\emptyset$ then
14:     $L\leftarrow L\cup\{(r,D_{r})\}$ $\triangleright$ explicit deny
15:   else
16:     $L\leftarrow L\cup\{(r,\emptyset)\}$ $\triangleright$ implicit deny
17:   end if
18:  end if
19:end for
20:return $L$
IV Dataset Collection and Request Generation
CloudFix works to repair cloud access control policies that are written in cloud access control languages. However there is currently no public dataset specifically curated for policy repair, which makes research in this area difficult. This is because policies model access control within organizations and if these policies were made public they would could create security vulnerabilities. To address this gap, we curate real-world AWS IAM policies from the AWS re:Post forum and make them available to the research community. An important component of policy repair is the availability of representative request sets, but obtaining such a dataset is non-trivial because AWS IAM request logs are also kept private. To overcome these gaps, we algorithmically generate synthetic requests directly from the structure of the policy by varying the proportions of must-allow and must-deny requests. Below, we discuss the details of dataset collection and request generation.
IV-A Policy Collection
The policy collection process involves three steps: first scraping policies from public sources, then detecting and categorizing valid JSON policies, and finally normalizing and repairing formatting inconsistencies. To collect a sufficiently large dataset of IAM policies, we sourced user posts from AWS re:Post. In this community forum, AWS users share questions and answers regarding AWS services and products. We specifically targeted posts containing IAM policies that either failed to function as intended or required correction. In total we scraped over 9,968 IAM-related posts. We then performed policy detection and categorization. Each post body was scanned for JSON objects resembling IAM policies. Detection required identifying bracketed structures containing keys such as Effect, Action, or Statement. All collected policies were then tested for syntactic correctness using Quacky [eiers2022quacky], a quantitative policy analyzer. Quacky revealed that many of the policies exhibited structural errors or formatting inconsistencies. To correct these issues, we implemented a naive syntactic repair tool to ensure we had a corpus of policies which adhered to standard IAM policy conventions. In total, we collected 282 faulty policies (syntactically correct but needing to be repaired) and 45 repaired policies (policies that were community-accepted as repairing the original broken policy). We summarize the statistics of our dataset in Table I.
TABLE I: Summary statistics of the AWS IAM policy dataset.
Metric Value
Total Policies 282
Total Statements 446
Avg. Statements per Policy 1.58
Min Statements per Policy 1
Max Statements per Policy 10
Unique Services 92
Unique Actions 664
Unique Resource Types 189
Cross-Service Policies 51
Effect Distribution
   Allow (Count / %) 400   (89.7%)
   Deny (Count / %) 44   (9.9%)
IV-B Request Generation
[…]
Algorithm 4 GenerateRequests
1:function GenerateRequests($\mathcal{P}$, $n$, $\rho$)
2:  $\mathcal{E}\leftarrow$ Extract($\mathcal{P}$) $\triangleright$ Extract policy elements
3:  $\mathcal{R}^{+}\leftarrow$ SampleAllowed($\mathcal{E}$, $\lfloor 0.6n\rfloor$)
4:  $\mathcal{R}^{-}\leftarrow$ SampleDenied($\mathcal{E}$, $n-|\mathcal{R}^{+}|$)
5:  $\mathcal{R}\leftarrow\mathcal{R}^{+}\cup\mathcal{R}^{-}$
6:  $\mathcal{M}\leftarrow$ RandomSubset($\mathcal{R}$, $\lfloor\rho\cdot n\rfloor$) $\triangleright$ Select subset to misclassify
7:  for each $r\in\mathcal{M}$ do
8:   $r.\text{Effect}\leftarrow$ Flip($r.\text{Effect}$) $\triangleright$ Change allow to deny or vice versa
9:   $r\leftarrow$ ModifyAttributes($r$) $\triangleright$ Change one of $(s,p,c)$
10:  end for
11:  return $\mathcal{R}$
12:end function
13:function SampleAllowed($\mathcal{E}$, $k$)
14:  $\mathcal{R}^{+}\leftarrow\emptyset$
15:  for $i\leftarrow 1$ to $k$ do
16:   $a\leftarrow$ Sample($\mathcal{E}.\text{actions}$)
17:   $s\leftarrow$ Sample($\mathcal{E}.\text{resources}$)
18:   $p\leftarrow$ Sample($\mathcal{E}.\text{principals}$)
19:   $c\leftarrow$ Sample($\mathcal{E}.\text{conditions}$)
20:   $\mathcal{R}^{+}\leftarrow\mathcal{R}^{+}\cup\{\langle a,s,p,c,\text{allow}\rangle\}$
21:  end for
22:  return $\mathcal{R}^{+}$
23:end function
24:function SampleDenied($\mathcal{E}$, $k$)
25:  $\mathcal{D}\leftarrow$ GenerateComplements($\mathcal{E}$)
26:  $\mathcal{R}^{-}\leftarrow\emptyset$
27:  for $i\leftarrow 1$ to $k$ do
28:   $(a,s,p,c)\leftarrow$ Sample($\mathcal{D}$)
29:   $\mathcal{R}^{-}\leftarrow\mathcal{R}^{-}\cup\{\langle a,s,p,c,\text{deny}\rangle\}$
30:  end for
31:  return $\mathcal{R}^{-}$
32:end function
V Experimental Design
The goal of our evalution is to answer the following research questions:
RQ1: Can CloudFix repair cloud access control policies?
RQ2: How effective is Fault Localization for policy repair?
RQ3: How does accuracy vary across LLMs with FL-guided Repair?
RQ4: How well do the repairs generated by CloudFix generalize to requests not seen during the repair process?
RQ5: Can CloudFix repair policies with requests generated from natural language descriptions of faulty policies?
V-A Experiment Setup
All experiments were run using a Rocky Linux 9.3 computer cluster (Blue Onyx). Each job was allocated a single NVIDIA L40S GPU with 46 GB of memory, two Intel(R) Xeon(R) Platinum 8562Y+ CPU cores, and 251 GB of system RAM. All experiments were executed with a 64 GB memory limit and a single GPU per task.
V-B Models
We evaluate our approach using four open-source instruction-tuned models of comparable size. Specifically, we used CodeLLaMA-7B-Instruct, a model trained for code synthesis and reasoning; Granite-3.3-8B-Instruct, an IBM model optimized for instruction following in software engineering tasks; DeepSeek-Coder-7B-Instruct, a model finetuned for code completion and reasoning, designed for software repair and program synthesis, and Llama3.2-3B-Instruct, an instruction-tuned text-only model that supports a maximum of 128K context window. We maintain similar prompts and model hyperparameters for all models used. All models were executed through Hugging Face’s transformers pipeline [wolf2019huggingface].
V-C Repair Approaches
We employ two repair approaches in our experiments. The first is a Baseline repair approach which uses a basic prompt to instruct the LLM to repair the policy. The Baseline receives the complete faulty policy, the requests that must be allowed and denied, and a generic repair instruction, where the model needs to identify faulty statements within the policy and repair them. The second is the Fault Localization guided approach, which is the crux of our framework and described in detail in Section III. We maintain the same user and system prompts for both approaches to ensure that the quality of the prompts does not confound the results.
Another important factor that influences repair accuracy is the size of the request. We use Algorithm 4 to generate 10, 20, 30, and 50 request sizes. To increase the complexity of the requests, we ensure that each request targets a unique resource–resource-action-condition-principal pair. Furthermore, our dataset exhibits a higher proportion of must-allow requests compared to must-deny requests, as shown in Table I. This distribution reflects realistic cloud usage patterns where most access requests are operations that must be allowed by access control policies.
V-D Evaluation Metric
Accuracy. We evaluate our approach using accuracy, which is defined as the percentage of request specifications that are correctly classified by the policy. $\text{Accuracy}=\frac{R_{c}}{R_{\text{total}}}\times 100\%$, where $R_{c}$ is the number of correctly classified requests and $R_{\text{total}}$ is the total number of requests. We define three accuracy categorys: 100% (complete repair), 80-99% (moderate repair), and <80% (failed repair).
Iteration, $I$. We evaluate the total number of iterations the repair takes to reach a complete, moderate, or failed repair. In our experiments, $I$ is set to 5.
Time. We measure the total repair duration, including the time that an LLM takes to generate a repair and the time the SMT solver takes to validate the repair across all iterations until either 100% is achieved or the max $I$ is reached.
V-E Experiments
RQ1: Can CloudFix repair cloud access control policies? Our first research question aims at testing whether our framework can automatically repair faulty policies such that all requests in the given specification are correctly classified.
Setup. To answer this question, we evaluate our framework on our dataset of 282 AWS IAM policies. For each policy, we generate must-allow and must-deny requests of 10, 20, 30, and 50 using Algorithm 4.
Results. Table III show that FL achieves 84.0% full repair rate on 10 requests compared to 48.2% for Baseline, and maintains 54.3% complete repair on 30 requests while Baseline drops to 22.3%. We see a performance degradation for 50 request sizes in both Baseline and Fault Localization, where the Baseline repairs 0 policies completely, while Fault Localization generates 7 completely repaired policies. The results demonstrate that FL consistently outperforms the Baseline as the request size increases, and gracefully degrades as Baseline performance dramatically decreases for higher request sizes.
We further performed a statistical significance analysis on the obtained results using a two-tailed t-test to compare the mean repair accuracies between the two approaches. The p-values below 0.001 for request sizes 10, 20, and 30 show strong statistical evidence (>99.9% confidence, see Table II, green cells) that FL genuinely outperforms Baseline. However, the p-value of 0.056 for request size 50 fails to meet the significance threshold of 0.05; meaning, the 3.85% improvement could be due to random chance rather than Fault Localization.
RQ2: How effective is Fault Localization for policy repair? Next, our second research question aims to determine whether FL improves repair accuracy. We compare our two approaches: (1) Baseline, which uses simple prompt consisting of only the faulty policy and requests, and (2) FL-guided repair, which additionally provides the fault localization information discussed in Section III.
Setup. To answer RQ2, we use the same experimental setup as RQ1, evaluating both approaches on the same 282 policies with request sizes of 10, 20, 30, and 50. Both approaches use similar LLM configuration, prompt structure, and repair iteration limits (we set $I$ to 5 iterations).
Results. Table III shows the distribution of repair accuracy across different request sizes. The FL-guided repair demonstrates significant improvements over the Baseline for request sizes 10, 20, 30, and 50. For request size 10, FL generates 237 complete repairs compared to only 136 for the Baseline. This advantage continues with FL achieving 170 complete repairs for 20 requests, 153 complete repairs for 30 requests, and 7 complete repairs for 50 requests.
For request size 10, the Baseline generates 101 policies in the <<80% category, while FL reduces this to only 19 policies. Similarly, for 30 requests, the Baseline generates 100 policies with <<80% accuracy compared to just 30 with FL. However, the Baseline achieves no 100% repairs but only 67 policies in the 80–99% range, while FL only achieves 7 complete repairs and produces 175 policies with <<80% accuracy compared to the 217 of the Baseline. We observed that as the request size increases, the repair task may exceed the LLM’s capability to integrate FL information effectively. We believe that the accuracy degradation at larger request sizes indicates that, while FL is highly effective for moderately complex repair tasks, additional techniques may be needed to maintain its benefits when the number of request sizes grows substantially.
TABLE II: Statistical Significance: Repair Accuracy of Baseline and FL
Request Size Baseline FL Difference p-value
(Accuracy %) (Accuracy %) ($\Delta$ $p$)
10 77.66 95.21 +17.55 <<0.001
20 83.40 90.69 +7.29 <<0.001
30 75.63 92.29 +16.67 <<0.001
50 62.44 66.28 +3.85 0.056
p << 0.001; not significant: p >> 0.05.
Figure 4: Average repair accuracy of the Baseline and Fault Localization (FL) approaches against the initial faulty policies across varying request set sizes.
TABLE III: Policy Repair Distribution by Request Size
Request Size
Category 10 20 30 50
Baseline Fault Localization Baseline Fault Localization Baseline Fault Localization Baseline Fault Localization
100% Fix 136 (48.22%) 237 (84.04%) 149 (52.83%) 170 (60.30%) 63 (22.34%) 153 (54.30%) 0 (0.00%) 7 (2.50%)
80–99% 45 (16.00%) 26 (9.22%) 62 (22.00%) 69 (24.50%) 119 (42.20%) 99 (35.10%) 67 (23.80%) 100 (35.46%)
<<80% 101 (35.82%) 19 (6.74%) 71 (25.20%) 43 (15.25%) 100 (35.50%) 30 (10.64%) 215 (76.24%) 175 (62.10%)
TABLE IV: Repair Time for Baseline and FL Across Request Sizes
Request Size
Metric 10 20 30 50
Baseline Fault Localization Baseline Fault Localization Baseline Fault Localization Baseline Fault Localization
Avg Iteration. 3.14 1.80 2.93 2.85 4.13 3.08 5.00 4.92
Avg Total Time Per Policy (s) 370.3 136.5 667.2 456.3 827.1 547.3 1778.6 1306.9
Avg SMT Time Per Policy (s) 2.98 0.27 8.58 0.44 11.05 1.11 8.26 3.92
Avg LLM Time Per Policy (s) 367.36 136.20 658.66 455.86 816.02 546.20 1770.29 1302.98
SMT Time % 0.80 0.20 1.30 0.10 1.30 0.20 0.50 0.30
LLM Time % 99.20 99.80 98.70 99.90 98.70 99.80 99.50 99.70
RQ3: How does accuracy vary across LLMs with FL-guided repair? Following our evaluation of repair accuracy across request sizes, we now explore the performance of LLMs in repairing access control policies.
Setup. We employed the instruction-tuned variants of CodeLlama-7B-Instruct, Llama-3.2-3B-Instruct, DeepSeek-Coder-7B-Instruct-v1.5, and Granite-3.3-8B-Instruct. We prompted all models using the Fault Localization-guided prompting strategy with a request size of 30 and for all of the 282 access control policies.
Results. Figure 5 shows the overall repair accuracy of each model. Notably, CodeLlama achieves the highest overall accuracy of 92.3%, followed by Llama of 63.7%. DeepSeek-Coder and Granite had a relatively similar overall repair accuracy of 54.0% and 53.7%. Table V further shows the accuracy breakdown. CodeLlama was able to repair 54.3% of the 282 policies. The low accuracy in the rest of the models is best explained in the third category, which shows an increasing number of failed repairs, going from left to right.
TABLE V: FL Repair Accuracy Comparison Across Different LLMs (Request Size 30)
Category Prompt CodeLlama Llama DeepSeek-Coder Granite
100% Fix FL-guided 153 (54.3%) 19 (6.7%) 37 (13.1%) 39 (13.8%)
80–99% FL-guided 99 (35.1%) 55 (19.5%) 32 (11.3%) 29 (10.3%)
<<80% FL-guided 30 (10.6%) 208 (73.8%) 213 (75.5%) 214 (75.9%)
RQ4: How well do the repairs generated by CloudFix generalize to requests not seen during the repair process? The goal of this experiment is to test whether our repair framework can identify and generalize to semantically similar access patterns rather than overfitting to specific requests. A motivating example for this experiment involves multiple requests targeting similar resources,
(s3:getObject,arn:aws:s3:::category/data-1)
(s3:getObject,arn:aws:s3:::category/data-2)
(s3:getObject,arn:aws:s3:::category/data-3)
which could be grouped together to allows access to
(s3:getObject,arn:aws:s3:::category/data-*)
Enumerating similar requests this way has undesired effect. It challenges the LLM’s ability to generalize. As a result, simpler models might try to patch the policy by adding a specific rule for each file. While this works to correctly classify the provided requests, it fails for semantically related other requests and requires the models to reason more. Therefore, we test whether LLMs can repair policies in a way that generalizes to semantically similar access requests.
Setup. To answer RQ4, we evaluate the generalizability of CloudFix. To enumerate the requests, we first pick one misclassified request from the request sets of each policy and enumerate its resource into 10 new variants, similar to those conforming to the pattern shown above. This generated a total of 20 requests. Next, for each policy, we generated the repair using both the Baseline and FL-guided approaches on these 20 requests. To evaluate the repair on unseen requests, we then synthesized 15 variant requests from the same requests and evaluated whether the repaired policies correctly classified these unseen but semantically similar requests.
Results. Figure 6 presents the repair and generalization accuracy for both approaches. The left panel shows that FL-guided repair achieves 82.8% (172 complete repairs) accuracy compared to the Baseline’s 79.9% (147 complete repairs), showing a 2.9% improvement over the initial 20.9% accuracy of the faulty policies. We can see from the right panel of the figure that the Baseline achieves 79.9% accuracy on unseen request variants, and FL achieves 84.9%. Interestingly, the Baseline generalization accuracy matches its repair accuracy (79.9%). This suggests that when the Baseline generates complete repairs, it produces repairs that generalize well to semantically similar requests. The FL approach shows a stronger generalization, achieving 84.9% on variants compared to its 82.8% repair accuracy.
Figure 5: An overall repair accuracy of four different LLMs using the FL-guided prompting strategy with a request size of 30.
Figure 6: The left panel compares the repair accuracy of the Baseline (79.9%) and FL-guided (82.8%) methods against the initial faulty policy (20.9%). The right panel evaluates how these repairs generalize to unseen requests, showing the FL-guided approach (84.9%) can generate more generalizable repairs than the Baseline (79.9%).
RQ5: Can CloudFix repair policies with requests generated from natural language descriptions of faulty policies Finally, we test whether CloudFix can completely repair policies using requests generated by LLMs.
Setup. To answer RQ5, we select 10 faulty policies from our dataset. As mentioned in Section IV, our AWS IAM dataset contains 45 policies that have faulty policy: the original, incorrect policy posted by a user experiencing access control issues, ground truth policy: the corrected policy provided by an AWS Cloud Engineer in the forum, and user intent: the natural language description of the request specifications or problem as described by the user. We first use GPT-5, accessed via it’s web interface, to synthesize requests. We follow the following procedure: for each policy, we prompted the model with the user intent, faulty policy, and ground truth policy and instruct it to generate 5 up to 10 valid requests. The criteria to generate a valid request were that they must be misclassified by the faulty policy, correctly classified by the ground truth, and encapsulate the user’s intent. We then used Algorithm 2 to validate that the requests were misclassified by the faulty policy and correctly classified by the ground truth policy. To validate if the requests align with the user’s expressed intent, we manually verified to ensure correctness. Next, we used CodeLlama-7B-Instruct to repair the 10 faulty policies with a maximum of 5 iterations and a target accuracy of 100%.
Results. Figure 7 shows that the Baseline approach achieves 85.2% repair accuracy, while FL achieves 86.4%, both showing improvements over the 46.7% initial accuracy. This demonstrates that our framework can be used to generate policy repairs using requests generated by LLMs from natural langauge descriptions of the users intents.
Figure 7: The results for CloudFix on requests generated from natural language descriptions. The marginal difference between the two methods suggests that the requests generated by the LLM from user intent were effective at encapsulating the nuanced semantic difference.
VI Discussion and Threat to Validity
In this study, our main contribution is to leverage LLMs for policy repair and showing that Fault Localization-guided prompt enables LLMs to repair with improved accuracy and time, particularly for moderately complex request sizes (RQ1, RQ2). Our results show that the choice of LLMs plays a key part in performance, as demonstrated in RQ3, with CodeLlama-7B-Instruct achieving the highest complete repair. Results obtained in RQ4 show that LLMs can be used to repair policies that generalize to unseen, semantically similar requests. We further showed that LLMs can be used to synthesize access requests, which can be used to repair faulty access control policies (RQ5). We noticed the baseline repair took significantly more time for all repairs than FL, as shown in Table IV. This was often because the Baseline was attempting to enumerate the requests in an attempt to repair the policy, which often resulted in generating a failed policy (<80%). Our study did not use a timeout because we planned to focus on fairly evaluating the repair capability of our framework. Finally, our study relied on access control policies we curated from AWS re:Post, and algorithmically generated requests based on the structure of these policies; however, these policies may not fully capture the nuanced access control policies in real-world. The generalizability of our findings is constrained by the size of requests. We observed a degradation of repair accuracy and an increment of the overall repair time as the number of requests went up to 50, where FL generated only 7 complete repairs. Future work should consider techniques to maintain better accuracy for higher request sizes.
Beyond repair accuracy of our framework, the practicality of CloudFix hinges on its computational efficiency. Our analysis in Table IV shows that the FL-guided approach is significantly time-efficient than the baseline counterpart. However, more than 99% of the repair time was LLM inference time; therefore, future work should investigate the trade-off between accuracy and efficiency in model selection.
VII Related Work
Access Control Policy Repair: Significant research has been conducted towards verifying and maintaining correct access control policies [10.5555/3161326, DetecingResolvingPolicyMisconfig, 10.1145/1978942.1979243, KERN2022103301, Jayaraman2014AutomatedAA, Son2013FixMU, UsabilityChallengesAccessControl, UsabilityConsiderationsAccessControl, EffectsPolicyConflictResoution]. Numerous access control policy repair techniques based on formal methods have been proposed to fix policies to adhere to the principle of least permissiveness. Prior work by [xu2016towards] introduces a fault localization and mutation-based policy repair method to debug and XACML access control policies. D’Antoni et al. [d2024automatically] proposed a synthesis-based framework to automatically reduce permissiveness in (AWS) IAM policies by leveraging access logs. Similarly, [peng2017towards] introduced fault localization and mutation-based policy repair techniques to produce a list of suspicious elements in XACML policies by correlating the test results and the test coverage information. They used mutation-based policy repair searches for fixes by mutating suspicious policy elements with predefined mutation operators. Eiers et al. [eiers2023quantitative] proposed a quantitative symbolic analysis framework to automatically repair overly permissive cloud access control policies. They encoded policy semantics as SMT formulas and used model counting to quantify and iteratively reduce the permissiveness of access control policies. However, their approach cannot repair semantically faulty policies.
LLM-Based Automated Program Repair: Recent methods have shown promising results that LLMs can be effectively applied to automated program repair [zhang2024systematic][li2025hybrid][hossain2024deep][xia2023automated][campos2025empirical]. These approaches frame bug fixing as a code generation problem. A recent work closely related to ours, [orvalho2025counterexample], introduces a formal method based on counterexamples to guide LLMs to provide more accurate program repairs. However, to the best of our knowledge, there is no prior work that has explored automated access control policy repair with LLMs guided by a formal method-based fault localization.
VIII Conclusion
In this work, we introduced CloudFix, the first automated framework to integrate LLMs with Formal Methods for access control policy repair. Manually repairing policies is error-prone and time-consuming. We address this problem by developing a framework to automatically create policy repairs with LLMs and by enhancing their reasoning capabilities with targeted repairs, enabled by Fault Localization. Our evaluation on a curated dataset of 282 real-world AWS policies showed that LLMs enhanced with Fault Localization can be used to repair cloud access control policies with significantly improved accuracy than the baseline approach.
IX Data Availability
The policy dataset supporting this paper’s findings can be found here: https://doi.org/10.5281/zenodo.17373778